# Bridging Perception and Planning: Towards End-to-End Planning for Signal Temporal Logic Tasks

### v1

@@ VI-B Implementation details
We train on a single NVIDIA A800 (80 GB) with batch size $12$ for $24$ epochs using AdamW (betas $0.9/0.95$) and three parameter groups: (i) weight-decayed ($10^{-2}$), (ii) norm/bias/embedding without decay, and (iii) router centroids (expert_key) with a $2\times$ larger learning rate and a small decay $10^{-3}$; the base learning rate is $1{\times}10^{-5}$. Unless noted, we set d=512d{=}512, horizon T=80T{=}80, and a $4$-layer decoder with $4$ heads and hidden width $128$; MoE is enabled in the last $2$ layers with B=6B{=}6 buckets, $E^{\prime}{=}2$ experts per bucket, temperature $\tau{=}1.0$, and in-bucket top-k=2k{=}2 (dropout $0.1$).
@@ VI-C Ablation Study
Table II reports ablations on three factors while keeping splits, schedule, and optimiser settings fixed: (a) the loss margin $\gamma$ in the STL hinge term of Eq. (12), (b) the sharpness $k$ of the smooth $\max/\min$ used in robustness aggregation, and (c) the image resolution/backbone.
@@ VI-D Main Results
We evaluate on five in-distribution (ID) subsets, single-F, single-G, only-AND, only-OR, only(AND-OR) and their union All (ID), and three aggregated out-of-distribution (OOD) suites (OOD-1/2/3 (All)). The primary metric is success rate, defined as jointly satisfying the STL specification, remaining collision-free, and meeting kinematic feasibility. All entries in Table I are percentages.
@@
The rollouts provide evidence that the model has learned image- and task-conditioned planning, producing collision-free and kinematically admissible trajectories across both ID and OOD layouts. Given only multi-iew image observations and finite-horizon STL specifications, the planner consistently satisfies the prescribed reach/sequence predicates while remaining within the drivable set and respecting per-step motion limits, without per-layout tuning. This indicates transfer of both scene understanding and symbolic goal structure rather than memorization of a particular map.
@@ VII Conclusion
In this work, we presented, to the best of our knowledge, the first end-to-end model S-MSP for STL-specified robotic control, mapping synchronized sensory inputs and symbolic specifications directly to feasible trajectories. A structure-aware sparse Mixture-of-Experts exploits STL compositionality and temporal scope, yielding faster convergence and high satisfaction and feasibility at low latency. For deployability, we augment the policy with a runtime repair layer that invokes an RRT-style local replanner to modify only violating subsegments while respecting dynamics and obstacles.
For training efficiency, the present work focus on STL fragments without nested temporal operators. In future work, we plan to address the challenge of nested operators within our framework. Additionally, we aim to integrate the policy with reinforcement learning for preference alignment and adaptive refinement under distributional shifts, enhancing robustness during deployment.

### v2

@@ VI-B Implementation details
We train on a single NVIDIA A800 (80 GB) with batch size $12$ for $24$ epochs using AdamW (betas $0.9/0.95$) and three parameter groups: (i) weight-decayed ($10^{-2}$), (ii) norm/bias/embedding without decay, and (iii) router centroids (expert_key) with a $2\times$ larger learning rate and a small decay $10^{-3}$; the base learning rate is $1{\times}10^{-5}$. Unless noted, we set d=512d{=}512, horizon T=80T{=}80, and a $4$-layer decoder with $4$ heads and hidden width $128$; MoE is enabled in the last $2$ layers with B=6B{=}6 buckets, $E^{\prime}{=}2$ experts per bucket, temperature $\tau{=}1.0$, and in-bucket top-k=2k{=}2 (dropout $0.1$). What’s more, the sampling probability $\beta$ in TSP is 0.5. The source codes are available at https://github.com/yc7421cmd/E2E-Signal-Temporal-Logic-Planner.git.
@@ VI-C Ablation Study
Table II reports All(ID) ablations on three factors while keeping splits, schedule, and optimiser settings fixed: (a) the loss margin $\gamma$ in the STL hinge term of Eq. (12), (b) the sharpness $k$ of the smooth $\max/\min$ used in robustness aggregation, and (c) the image resolution/backbone.
@@ VI-D Main Results
We evaluate on five in-distribution subsets, single-F, single-G, only-AND, only-OR, only(AND-OR) and their union set All (ID), and three aggregated out-of-distribution (OOD) suites (OOD-1/2/3 (All)). The primary metric is success rate, defined as jointly satisfying the STL specification, remaining collision-free, and meeting kinematic feasibility. All entries in Table I are percentages.
(a) ID
(b) OOD-1
(c) OOD-2
(d) OOD-3
@@
The rollouts provide evidence that the model has learned image- and task-conditioned planning, producing collision-free and kinematically admissible trajectories across both ID and OOD layouts. Given only multi-iew image observations and finite-horizon STL specifications, the planner consistently satisfies the prescribed reach/sequence predicates while remaining within the drivable set and respecting per-step motion limits, without per-layout tuning. This indicates transfer of both scene understanding and symbolic goal structure rather than memorization of a fixed particular map.
@@ VII Conclusion
In this work, we presented, to the best of our knowledge, the first end-to-end model S-MSP for STL-specified robotic control, mapping synchronized sensory inputs and symbolic specifications directly to feasible trajectories. A structure-aware sparse Mixture-of-Experts exploits STL compositionality and temporal scope, yielding faster convergence and high satisfaction and feasibility at low latency. For deployability, we augment the policy with a runtime repair layer that invokes an RRT-style local replanner to modify only violating subsegments while respecting dynamics and obstacles. The experimental results validate the effectiveness and robustness of the proposed approach.
In this work, we restrict attention to STL fragments without nested temporal operators. Future work will extend the framework to nested operators and incorporate reinforcement learning for preference alignment and adaptation under distributional shifts to improve deployment robustness.